OpenStudy (anonymous):

If g(x) = 2x- x + 5, then g (2)=

OpenStudy (lgbasallote):

sub 2 into every x of g(x)

OpenStudy (anonymous):

just sub 2 for x

OpenStudy (anonymous):

is it 7

OpenStudy (anonymous):

just plug in 2 for every X value g(2) = 2(2) - 2 + 5 = 4 - 2 +5 = -2 + 5 = 3

OpenStudy (anonymous):

sorry 2 + 5 not -2 + 5

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

whats the final answer is it 3 or 7

OpenStudy (anonymous):

its 7

OpenStudy (anonymous):

ok thanks